OpenStudy (anonymous):

write the slope intercept form of the line (-5, 6) (5, 2)

OpenStudy (anonymous):

slope is change in y over change in x, or (y1 - y0) / (x1 - x0). (2 - 6)/(5 - -5) = -4/10 = -2/5. That means for every 5 you move in the positive x direction, the line drops 2 in the y y direction. So moving from (-5,6) forward 5 in the x direction, the line will intersect the y axis at (0,4). Now you can put this in the form of y = mx + b, where m is slope and b is y-intercept.

OpenStudy (anonymous):

http://bit.ly/lUjHeU

OpenStudy (anonymous):

I found the slope to be -2/5, but still confused